Solve for x:

log base 4 of(x^2+3x) - log base 4 of(x+5)=1

Answer by stanbon(75887)

log base 4 of(x^2+3x) - log base 4 of(x+5) = 1
-------------
log4[(x^2+3x)/(x+5)] = 1
----
(x^2+3x)/(x+5) = 4
---
x^2+3x = 4x+20
x^2-x-20 = 0
Factor:
(x-5)(x+4) = 0
x = 5 or x = -4
=======================
Cheers,
Stan H.
==============